# DLF: Extreme Image Compression with Dual-generative Latent Fusion

### v1

@@ 3.3 Two-stage Training Strategy
In the latent alignment stage, we apply a rate-distortion loss in the latent space for alignment. Given the pretrained pixel generator, we use its corresponding encoder [12] (i.e. the auxiliary encoder in Fig.7) to generate $\tilde{h}$ to supervise the latent reconstruction $\hat{h}$. The rate is calculated based on the distribution of the quantized detail latents. After the first stage, we further fine-tune the entire model using the pixel-space generative loss [12]. During the training process, the detail branch will learn the spatial-adaptive compression and allocate more bits to the most important regions for fidelity. More details about the block design can be found in our supplementary material.
@@ 4.1 Implementation Details
Training details. DLF is trained on the subset of the OpenImages train set [19]. During stages I, we train the model on randomly cropped $256\times 256$ patches with a batch size of 8, using a fixed bitrate weight $\lambda$ of $24.0$. In this stage, we incorporate the pretrained 1-D tokenizer [47] and freeze its weights. This offers a strong initialization to leverage its semantic prior, facilitating faster convergence. For stage II, we fine-tune the model using randomly cropped $512\times 512$ patches with a batch size of 4. Here $\lambda$ is set to $\{5.8,8.5,16.0,28.0\}$ to achieve different bitrates.
@@
Figure 10: Qualitative examples on the MS-COCO 30K dataset.
@@
Comparison methods. We compare our DLF with the traditional codec VVC [6], the MSE-optimized neural codec TCM [27], and various generative codecs. They include the VAE-based HiFiC [32], MS-ILLM [35], the tokenizer-based GLC [16], the dual-branch-based HybridFlow [30] and the diffusion-based PerCo [7], DiffEIC [24] and RDEIC [25]. We evaluate generative methods exclusively on the MS-COCO 30K dataset, as it is primarily used for assessing generation quality. Diffusion-based methods are evaluated on the cropped CLIC2020 dataset instead of full resolution due to their limitations in high-resolution compression.
@@ 4.2 Main Results
Quantitative Evaluation. In Fig. 8, we present the rate-distortion curves across various datasets. For fidelity metrics, DLF achieves superior performance than all other methods, yielding significantly lower (better) LPIPS and DISTS scores at the same bitrates. We also compute the BD-Rate [4] on LPIPS and DISTS in Table 1. Using MS-ILLM as anchor, DLF achieves bitrate savings of 43.05% for LPIPS and 67.82% for DISTS on the Kodak dataset, significantly outperforming the recent tokenizer-based GLC and diffusion-based DiffEIC. Additionally, our compression ratio exceeds HybridFlow by more than 2.6×2.6$\times$2.6 ×, demonstrating a substantial improvement in dual-branch compression. Regarding generation realism measured by FID, our method outperforms all non-diffusion-based approaches on the MS-COCO 30K dataset while showing competitive results against advanced diffusion-based method PerCo. Surprisingly, on the cropped CLIC2020 dataset at $768\times 768$ resolution, DLF surpasses PerCo in terms of FID by a large margin, underscoring its superiority in preserving realism on high-quality datasets. As shown in Fig. 11, DLF achieves the best rate-distortion performance for both FID and DISTS on cropped CLIC2020. These results demonstrate the effectiveness of DLF, highlighting the great potential of our interactive dual-branch coding design.
Qualitative Evaluation. In Fig. 9, we present a visual comparison from the CLIC2020 test set. The MSE-optimized TCM results in severely blurry reconstructions. Similarly, the generative codecs HiFiC and MS-ILLM struggle to generate realistic details at a low bitrate around 0.01 bpp, resulting in blurred details and additional artifacts. By contrast, DLF produces very clear and high quality reconstruction with excellent fidelity. Additional examples from the MS-COCO 30K dataset can be found in Fig. 10. In this scenario, MS-ILLM continues to exhibit blurry reconstructions. Although diffusion-based methods PerCo and DiffEIC achieve greater realism, they also introduce content distortions (e.g., on the grape) in their reconstructions. Additionally, the reconstructions of PerCo are accompanied by color deviation, negatively affecting visual quality. Compared to these methods, DLF produces visually appealing reconstructions with fewer artifacts, accurate colors, and higher fidelity details, even at lower bitrates. Refer to our supplementary material for more visual results.
@@ 4.3 Ablation Study
In this section, we conduct ablation studies to evaluate the impact of the dual-branch coding architecture, the cross-branch interactive design, and scalar-quantization (SQ) based detail coding design. All model variations are evaluated using the Kodak and CLIC2020 datasets. BD-Rate results are presented in Table 2.
Dual-branch coding architecture. We remove the entire detail branch, including all IT modules (denoted as w/o detail), and compress the image using only the semantic branch. In this variation, the bitrate is adjusted solely by varying the number of VQ tokens. As shown in Table 2, removing detail branch brings a loss in BD-Rate over 47% on the CLIC2020 dataset, as measured by LPIPS and DISTS. These findings validates that the dual-branch coding is superior than single-branch semantic coding.
Cross-branch Interactive Design. In this variation, we remove all IT modules, eliminating interactions between the two branches (denoted as w/o interactive). This setting results in the most severe performance drop across all metrics and datasets, with bitrate loss exceeding 60%. It indicates that using two parallel but independent branches leads to substantial redundancy between the two bottlenecks, underscoring the crucial role of our interactive design.
SQ-based Detail Quantization. In this variation, we replace the scabilityntization in the detail branch with vanilla vector quantization (denoted as w/ VQ detail). Huffman coding is applied to encode VQ indices, and the bitrate is adjusted via the codebook size. It also show a non-trival performance drop, indicating that the limited quantization space of the VQ codebook is insufficient for capturing various individual-level details, thus highlighting the effectiveness of our spatially-adaptive SQ design. Furthermore, the dual-VQ bottleneck does not outperform the single branch solution, indicating that our model requires a hybrid quantization strategy of VQ-SQ to achieve the best performance.
@@ 4.4 Complexity Analysis
We compare the average coding time (with standard deviation) in In Table 3. Compared to MS-ILLM, DLF incorporates a larger model thus has a relatively slower coding speed. However, the larger model enables stronger generative capabilities, which is the basis of high-quality reconstruction at extremely low bitrates of 0.01 bpp. Without sufficient generative capabilities, MS-ILLM produces quite blurry reconstructions. Some recent extreme image codecs have introduced large diffusion models (e.g., PerCo and DiffEIC). Notably, DLF offers a much faster decoding speed than diffusion-based methods with improved coding performance, which is a significant progress in adopting large models for extreme image compression.
@@ A.1 Stage 1: Latent Alignment
To ensure that the detail branch captures the most significant individual-level contents while discarding minor ones, we impose a rate constraint on the detail features. Consequently, we design the latent domain rate-distortion loss as follows:
$\mathcal{L}_{\text{Stage-1}}=||\hat{h}-\tilde{h}||^{2}_{2}+\lambda\cdot\mathcal{R}(\hat{y}_{d})$ (6)
where $\hat{h}$ is the decoded latent generated by the latent adaptor, and the supervision latent $\tilde{h}$ is obtained from a pretrained auxiliary encoder (i.e., the VQGAN encoder [12]). Additionally, $\mathcal{R}(\hat{y}_{d})$ represents the bitrates of the quantized detail latent $\hat{y}_{d}$. The weight $\lambda$ controls the trade-off between rate and distortion items.
At the start of this stage, we load the pretrained weights of the 1-D tokenizer [47] into the semantic branch. These weights are fixed during this stage to maintain their initialization benefits. This approach provides a strong initialization, accelerating the training process and aiding faster convergence. In our model, the semantic codebook size is set to 4096.
@@ A.2 Stage 2: End-to-end Fine-tune
In this stage, we fine-tune the entire model, including all parameters in both branches, the latent adaptor, and the pixel generator. To achieve superior reconstruction quality, we employ the pixel-domain rate-distortion loss, which is defined as:
@@
$\mathcal{L}_{\text{pixel}}=||x-\hat{x}||+\mathcal{L}_{\text{LPIPS}}(x,\hat{x})+\lambda_{\text{adv}}\cdot\mathcal{L}_{\text{adv}}(x,\hat{x})$ (8)
$\mathcal{L}_{\text{codebook}}=||\text{sg}(y_{s})-\hat{y_{s}}||+\beta\cdot||\text{sg}(\hat{y_{s}})-y_{s}||$ (9)
@@ Appendix B Model Architectures
We illustrate the model architecture with the detailed hyper-parameters. Fig. 12 shows the architecture of the semantic tranform, the detail transform and the latent adaptor. Fig. 13 illustrates the downsample module before the scalar quantization and the upsample module after the quantization within the detail branch. Lastly, Fig. 15 depicts the design of the Interactive Transform (IT) module.
@@ C.2 Quantitative Results
We provide the additional KID [3] results on the CLIC2020 and the MS-COCO 30K datasets on the Fig 14. We present the evaluation results on the DIV2K dataset [1] at full resolution in Fig. 16. Large version of the Rate-distortion curves is shown in Fig. 22. Diffusion-based methods [7, 24] are excluded from evaluation on this dataset due to their large memory requirements, which exceed the capacity of our GPU (A100 with 40GB memory). From Fig. 16, we can see that our method outperforms MS-ILLM [35] across all reference and no-reference perceptual metrics, demonstrating its effectiveness. Additionally, we evaluate traditional pixel-level distortion metrics, PSNR and MS-SSIM [42], for a more comprehensive analysis, as shown in Fig. 17. It is worth noting that at extreme low bitrates, pixel-level distortion becomes too severe (e.g., the PSNR of VTM drops below 25 dB), making these metrics less meaningful for evaluating visual quality. Although our model does not exhibit the best pixel-level distortion metrics, it still provides the most visually appealing reconstructions with high fidelity, as demonstrated in previous sections.
@@ C.3 Qualitative Results
In this section, we provide more visual examples across the CLIC2020 test [40] (Fig. 18, 19, 20) and MS-COCO 30K [26] (Fig. 21) datasets. From these examples, we can find that DLF achieves the best quality with the lowest bitrate.
Figure 18: Qualitative examples on the CLIC2020 dataset (full resulution).
Figure 19: Qualitative examples on the CLIC2020 dataset (full resulution).
Figure 20: Qualitative examples on the CLIC2020 dataset (768$\times$768).

### v3

@@ 3.3 Two-stage Training Strategy
In the latent alignment stage, we apply a rate-distortion loss in the latent space for alignment. Given the pretrained pixel generator, we use its corresponding encoder [13] (i.e. the auxiliary encoder in Fig.7) to generate $\tilde{h}$ to supervise the latent reconstruction $\hat{h}$. The bitrate is calculated based on the distribution of the quantized detail latents. After the first stage, we further fine-tune the entire model using the pixel-space generative loss [13]. During the training process, the detail branch will learn the spatial adaptive compression and allocate more bits to the most important regions for fidelity. Further details on the training and block design are provided in the supplementary material.
@@ 4.1 Implementation Details
Training details. DLF is trained on the subset of the OpenImages train set [1]. During stage I, we train the model on randomly cropped $256\times 256$ patches with a batch size of 8, using a fixed bitrate weight $\lambda$ of $24.0$. In this stage, we incorporate the pretrained 1-D tokenizer [53] and freeze its weights. This offers a strong initialization to leverage its semantic prior, facilitating faster convergence. For stage II, we fine-tune the model using randomly cropped $512\times 512$ patches with a batch size of 4. Here $\lambda$ is set to $\{5.8,8.5,16.0,28.0\}$ to achieve different bitrates.
@@
Figure 10: Qualitative examples on the MS-COCO 30K dataset. More comparisons are in supplementary.
@@
Comparison methods. We compare our DLF with the traditional codec VVC [7], the MSE-optimized neural codec TCM [29], and generative codecs. They include the VAE-based HiFiC [34], MS-ILLM [36], the tokenizer-based GLC [17], the dual-branch-based HybridFlow [32] and the diffusion-based PerCo [21], DiffEIC [26] and RDEIC [27]. We evaluate generative methods exclusively on the MS-COCO 30K dataset, as it is primarily used for assessing generation quality. Diffusion-based methods are evaluated on the cropped CLIC2020 dataset instead of full resolution due to their limitations in high-resolution compression.
@@ 4.2 Main Results
Quantitative Evaluation. In Fig. 8, we present the rate-distortion curves across various datasets. For fidelity metrics, DLF achieves superior performance than all other methods, yielding considerably lower (better) LPIPS and DISTS scores at the same bitrates. We also compute the BD-Rate [5] on LPIPS and DISTS in Table 1. Using MS-ILLM as anchor, DLF achieves bitrate savings of 43.05% for LPIPS and 67.82% for DISTS on the Kodak dataset, significantly outperforming the recent tokenizer-based GLC and diffusion-based DiffEIC. Additionally, our compression ratio exceeds HybridFlow by more than $2.6\times$, demonstrating a substantial improvement in dual-branch compression. Regarding generation realism measured by FID, our method outperforms all non-diffusion-based approaches on the MS-COCO 30K dataset while showing competitive results against advanced diffusion-based method PerCo. Surprisingly, on the cropped CLIC2020 dataset at $768\times 768$ resolution, DLF surpasses PerCo in terms of FID by a large margin, underscoring its superiority in preserving realism on high-quality datasets. As shown in Fig. 11, DLF achieves the best rate-distortion performance for both FID and DISTS on cropped CLIC2020. These results demonstrate the effectiveness of DLF, highlighting the great potential of our interactive dual-branch coding design.
Qualitative Evaluation. In Fig. 9, we present a visual comparison from the CLIC2020 test set at a resolution of $768\times 768$. The VAE-based MS-ILLM struggle to generate realistic details at a low bitrate, resulting in blurred details and additional artifacts. While diffusion-based methods PerCo and DiffEIC improve realism, they introduce noticeable content distortions (e.g., in vegetation). Additional examples from the MS-COCO 30K dataset (Fig. 10) show similar trends: MS-ILLM remains blurry, and PerCo and DiffEIC fail to maintain fidelity (e.g., the distorted grape). Compared to these methods, DLF reconstructs visually appealing images with fewer artifacts, accurate colors, and finer details, even at lower bitrates. More visual results are available in the supplementary material.
@@ 4.3 Ablation Study
In this section, we conduct ablation studies to evaluate the impact of the dual-branch coding architecture, the cross-branch interactive design, and the scalar quantization (SQ) based detail coding design. All model variants are evaluated using the Kodak and CLIC2020 datasets. BD-Rate results are presented in Table 2.
Dual-branch coding architecture. We remove the entire detail branch, including all IT modules (denoted as w/o detail), and compress the image using only the semantic branch. In this variant, the bitrate is adjusted solely by varying the number of VQ tokens. As shown in Table 2, removing detail branch brings a loss in BD-Rate over 47% on the CLIC2020 dataset, as measured by LPIPS and DISTS. These findings validate that the dual-branch coding is superior than single-branch semantic coding.
Cross-branch Interactive Design. In this variant, we remove all IT modules, eliminating interactions between the two branches (denoted as w/o interactive). This setting results in the most severe performance drop across all metrics and datasets, with bitrate loss exceeding 60%. It indicates that using two parallel but independent branches leads to substantial redundancy between the two bottlenecks, underscoring the crucial role of our interactive design.
SQ-based Detail Quantization. In this variant, we replace the scalar quantization in the detail branch with vanilla vector quantization (denoted as w/ VQ detail). Huffman coding is applied to encode VQ indices, and the bitrate is adjusted via the codebook size. It also shows a non-trivial performance drop, indicating that the limited quantization space of the VQ codebook is insufficient for capturing various individual-level details, thus highlighting the effectiveness of our spatially adaptive SQ design. Furthermore, the dual-VQ bottleneck does not outperform the single branch solution, indicating that our model requires a hybrid quantization strategy of VQ-SQ to achieve the best performance.
@@ 4.4 Complexity Analysis
We compare the average coding time (with standard deviation) in Table 3. Compared to MS-ILLM, DLF incorporates a larger model and thus has a relatively slower coding speed. However, the larger model enables stronger generative capabilities, which form the basis of high-quality reconstruction at extremely low bitrates of 0.01 bpp. Without sufficient generative capabilities, MS-ILLM produces noticeably blurry reconstructions. While compared to recent extreme image codecs that incorporate large diffusion models (e.g., PerCo and DiffEIC), DLF offers a much faster decoding speed than those methods, along with improved coding performance—representing considerable potential in adopting large tokenizer models for extreme image compression.
@@ A.1 Stage 1: Latent Alignment
To ensure that the detail branch captures the most significant contents while discarding minor ones, we impose a rate constraint on the detail features. Consequently, we design the latent domain rate-distortion loss as follows:
[…]
where $\hat{h}$ is the decoded latent generated by the latent adaptor, and the supervision latent $\tilde{h}$ is obtained from a pretrained auxiliary encoder (i.e., the VQGAN encoder [13]). Additionally, $\mathcal{R}(\hat{y}_{d})$ denotes the bitrate of the quantized detail latent $\hat{y}_{d}$, estimated using the quadtree-partition-based spatial context module [24]. The weight $\lambda$ controls the trade-off between rate and distortion items.
At the start of this stage, we load the pretrained weights of the 1-D tokenizer [53] into the semantic branch. These weights are fixed during this stage to maintain their initialization benefits. This approach provides a strong initialization, accelerating the training process and aiding faster convergence. The semantic codebook size is set to 4096.
@@ A.2 Stage 2: End-to-end Fine-tune
In this stage, we fine-tune the entire model, including all parameters in both branches, the latent adaptor, and the pixel generator. To achieve superior reconstruction quality, we employ the pixel-domain rate-distortion loss:
@@
[…]
@@ Appendix B Model Architectures
We illustrate the model architecture with the detailed hyper-parameters. Fig. 12 shows the architecture of the semantic transform, the detail transform and the latent adaptor. Fig. 13 illustrates the downsample module before the scalar quantization and the upsample module after the quantization within the detail branch. Lastly, Fig. 15 depicts the design of the Interactive Transform (IT) module.
@@ C.2 Quantitative Results
We provide the additional KID [4] results on the CLIC2020 and the MS-COCO 30K datasets on the Fig. 14. We present the evaluation results on the DIV2K dataset [2] at full resolution in Fig. 16. Large version of the rate-distortion curves is shown in Fig. 23. Diffusion-based methods [8, 26] are excluded from evaluation on this dataset due to their large memory requirements, which exceed the capacity of our GPU (A100 with 40GB memory). From Fig. 16, we can see that our method outperforms MS-ILLM [36] across all reference and no-reference perceptual metrics, demonstrating its effectiveness. Additionally, we evaluate traditional pixel-level distortion metrics, PSNR and MS-SSIM [45], for a more comprehensive analysis, as shown in Fig. 17. It is worth noting that at extreme low bitrates, pixel-level distortion becomes too severe (e.g., the PSNR of VTM drops below 25 dB), making these metrics less meaningful for evaluating visual quality. Although our model does not exhibit the best pixel-level distortion metrics, it still provides the most visually appealing reconstructions with high fidelity, as demonstrated in previous sections.
@@ C.3 Qualitative Results
In this section, we provide more visual examples across the CLIC2020 [42] (Full resolution: Fig. 18, 19, 20; $768\times 768$: Fig. 21) and MS-COCO 30K [28] (Fig. 22) datasets. From these examples, we can find that DLF achieves the best quality with the lowest bitrate cost.
Figure 18: Qualitative examples on the CLIC2020 dataset (full resolution).
Figure 19: Qualitative examples on the CLIC2020 dataset (full resolution).
Figure 20: Qualitative examples on the CLIC2020 dataset (full resolution).
Figure 21: Qualitative examples on the CLIC2020 dataset ($\times768$).